# PepTune: De Novo Generation of Therapeutic Peptides with Multi-Objective-Guided Discrete Diffusion

### v1

@@ 3 Discussion
In this work, we introduce PepTune, a generative framework that achieves multi-objective optimization directly in discrete sequence space. By leveraging MCTS for guidance, PepTune identifies Pareto-optimal peptide SMILES sequences conditioned on diverse therapeutic properties such as binding affinity, solubility, membrane permeability, and hemolysis. Unlike previous guidance methods, which struggle with gradient estimation [27] or rely on projections between continuous and discrete spaces, PepTune operates natively in the discrete latent space. Our approach combines exploration through batched unmasking and reward-based exploitation of classifier predictions, ensuring valid peptide structures with a state-dependent masking schedule and a straight-through estimator-based penalty. Most importantly, unlike recent binder design methods [12, 14, 15, 87], PepTune requires no obligate target three-dimensional structures or predicted structures (only the target sequence), enabling peptide design to conformationally diverse proteins, optimized for properties beyond local geometric interactions.
Despite its strengths, PepTune relies on synthetic peptide data (e.g., CycloPs [88]) and rare non-natural amino acids (nAAs) [40], which may increase synthesis complexity and costs. While we address this through feature-rich embeddings from a pre-trained chemical language model [45, 89], improving high-quality labeled datasets remains critical for enhancing property prediction accuracy. Furthermore, while we evaluate binding using an external, state-of-the-art docking strategy via AutoDock Vina [48], there is a lack of biophysical models for other properties optimized via PepTune, including peptide solubility, hemolysis, and membrane permeability, outside of existing predictor algorithms. As such, we are currently conducting in vitro assays to confirm these properties of our generated peptides, and will update the manuscript as these results are obtained.
@@ 4 Methods
Dynamic Batching. We applied dynamic batching to handle variable-length token sequences and increase computational efficiency. Inspired by ESM-2’s dynamic batching technique [44], input SMILES are sorted by length to maximize the utility of GPU memory. The maximum token size is 16k per GPU.
SMILES Tokenization To enable the novel generation of non-natural amino acids containing cyclizations and diverse backbone and side-chain modifications, we trained our generative diffusion model on Simplified Molecular-Input Line-Entry System (SMILES) [32] representations of peptides. We experimented with several tokenization schemes that capture common motifs in the training data to enhance the generation of valid peptide SMILES. We find that the SMILES Pair Encoding (SPE) tokenization scheme [95] with the PeptideCLM [45] vocabulary of 581 SMILES tokens and 5 special tokens with an average length of 4 characters per token, demonstrated superior performance, generating precise but valid peptides (Appendix D.2).
Classifier Data and Training. We trained our membrane permeability XGBoost regression model using 34,853 experimentally-validated peptide SMILES, with 22,040 SMILES sequences obtained from the ChEMBL database [98] and 7451 sequences from the CycPeptMPDB database [94].
We collected binding affinity, solubility, hemolysis, and non-fouling data for the classifier training [39, 30]. The binding affinity data was split based on binding affinity distributions with a 0.8/0.2 ratio. The classifier was trained with cross-attention between ESM-2 protein embeddings and PeptideCLM SMILES representations [47]. The hyperparameters were chosen with 50 trials of OPTUNA search [99]. For other classifiers, data were randomly shuffled and split into 0.8/0.1/0.1 ratio for train, validation, and test. XGBoost classifiers [100, 101] were applied on PeptideCLM embeddings with 50 trials of OPTUNA search for the optimal tree hyperparameters.
@@ 4.2 Unconditional Masked Discrete Diffusion Model
Notation. Let $\mathbf{x}_{0}\in\{0,1\}^{|\mathcal{V}|}$ represent the one-hot vector of a token in a sequence in the training data and $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\in\Delta^{|\mathcal{V}|}$ be the vector of predicted token probabilities across the vocabulary $\mathcal{V}$ given the current state $\mathbf{z}_{t}$ at time $t$. In most contexts, $\mathbf{x}_{0}$ will be used to denote a single token, but when discussing the full sequence, $\mathbf{x}_{0}^{(\ell)}$ is used to denote the token at position $\ell$ in the sequence. Let $T$ denote the total number of time steps in the discrete forward and reverse diffusion processes and $t(n)\in[0,1]$ denote a time step in the forward and backward diffusion process and let $s(n)=t(n)-\frac{1}{T}$ denote the previous time step with $n\in[1\dots T-1]$. Then, let $\mathbf{z}_{t(n)}$ and $\mathbf{z}_{s(n)}$ denote the state of a specific token at time $t(n)$ and $s(n)$ in the diffusion process, respectively. In some cases, $t(n)$ is simplified to $t$. Let $\alpha_{t}(\mathbf{x}_{0}):\mathbf{R}^{|\mathcal{V}|}\to\mathbb{R}$ denotes a function that takes a token $\mathbf{x}_{0}$ and outputs a value in [0,1] representing the probability of remaining unmasked at time $t$ in the forward diffusion process. Let $\mathbf{b}\in\mathbb{R}^{|\mathcal{V}|}$ denote a vector with ones at indices of peptide bond tokens.
State-Dependent Masking Schedule. Since all peptides follow a distinct SMILES structure consisting of un-modified or modified peptide bonds before and after each central carbon atom with an amino acid side chain, we hypothesized that applying state-dependent masking and unmasking schedules would allow the reverse diffusion process to learn to unmask the crucial structural components of a peptide SMILES that are common across all peptides before filling in the segments between with diverse amino acid side-chains.
Therefore, we devised a masking schedule where the probability of masking a token within a peptide bond increases at a slower rate in earlier times $t$ in the masking process compared to non-peptide bond tokens. To achieve this, we define the discrete-time log-linear masking schedule $\sigma(t)=\log(1-t)$ for non-peptide bond tokens and the log-polynomial masking schedule $\sigma(t)=-\log(1-t^{w})$ for peptide-bond tokens. We show in Appendix B.1 that the continuous-time probability of remaining unmasked at time $t$ in the forward diffusion process is given by the function $\alpha_{t}(\mathbf{x}_{0}):\mathbb{R}^{|\mathcal{V}|}\to\mathbb{R}$ that takes the vector encoding the token $\mathbf{x}_{0}$ and returns a probability.
@@
where $\mathbf{b}$ represents the set of one-hot encoding vectors for tokens that make up all types of peptide bonds, including all modified peptide-bond tokens. Since the probability of transitioning to a [MASK] token at time $t$ is given by $1-\alpha_{t}(\mathbf{x}_{0})$, there is a lower probability $t^{w}$ for $t\in[0,1]$ of masking a peptide bond token than the probability $t$ of masking a non-peptide bond token, especially in earlier time steps for smaller $t$ (Fig. 15A). As $t\to 1$, $\alpha_{t}(\mathbf{x}_{0})\to 0$ for both peptide and non-peptide bond tokens, ensuring that the model can learn to reconstruct the full token sequence during the reverse diffusion process.
Therefore, we define the state-dependent forward transition matrix as
@@
To estimate the reverse posterior, we define a parameterized model $\mathbf{x}_{\theta}(\mathbf{z}_{t},t):\mathcal{V}\times[0,1]\to\Delta^{N}$ that takes the masked sample at time $t$ and predicts a vector of token probabilities for the clean sample. Substituting $\mathbf{x}=\mathbf{x}_{\theta}(\mathbf{z}_{t},t)$ gives us the following reverse diffusion transition distribution:
@@
SUBS Parameterization. Following Sahoo et al. [22], we parameterize the reverse diffusion model using SUBS parameterization, which enforces zero-masking probability and carry-over unmasking. This strategy enforces the constraints applied in the forward diffusion process and has been shown to minimize test perplexity [103].
@@
    Zero Masking Probability. The forward process operates under the assumption that a token can only be masked once across times $t=0\to 1$. It follows that the probability of a sequence being masked in the reverse diffusion process is 0, we set the probability of the sequence being masked in the reverse diffusion model to be negative infinity such that if a token is unmasked at time $t$, it remains unmasked for all timesteps from $t\to 0$.
@@
    Carry-Over Unmasking. For each transition in the forward pass, all tokens either remain unchanged or are masked. Therefore, in the reverse process, when a token is unmasked at time $t$, the unmasked token is carried over all time steps from $t\to 0$.
    $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)=\mathbf{z}_{t}$ (10)
@@ 4.3 Loss Functions
Training on samples masked for continuous values of $t\sim\mathcal{U}(0,1)$ yields a tighter lower bound compared to discrete values of $t$. When the predicted probability distribution $\mathbf{x}_{\theta}(\mathbf{z}^{\ell}_{t},t)$ is exactly the one-hot encoding vector $\mathbf{x}_{0}^{\ell}$ for each token at position $\ell$ in the true sequence, the loss reduces to 0, which supports our objective.
@@
Given that peptide bonds form the fundamental backbone structure of a peptide, our state-dependent masking strategy for peptide bonds acts as a peptide bond loss that introduces a higher penalty when the token predictions at positions of peptide bonds are inconsistent from the ground truth tokens during training, forcing the model to learn the fundamental structure of peptide SMILES strings in a vast space of non-peptide SMILES strings.
Invalid Peptide Loss. To further discourage the generation from predicting token logits that produce invalid peptide SMILES, we incorporate a loss to penalize sampling of invalid peptide SMILES during training by taking the argmax of the predicted logits and passing the sampled SMILES string into a filter that classifies it as a valid or invalid peptide. However, since the argmax function is not differentiable, we use the softmax probability of the sampled token to scale the penalty score which acts as a scalar multiplier in the loss function.
Given a sequence of one-hot encoding vectors $\tilde{\mathbf{x}}^{(\ell)}\in\{0,1\}^{K}$ of the tokens with the highest probability from the predicted logits $\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)$, we aim to minimize the penalty generated by a validity function that returns 0 when the argmax sampled sequence is a valid peptide SMILES string and 1 when the sequence does not translate to a peptide. To ensure differentiability, we define the loss function as the penalty multiplied by the softmax probability sampling the given token in an invalid sequence.
$\mathcal{L}_{\text{invalid}}$ $=\sum_{\ell=1}^{L}\tilde{\mathbf{x}}_{0}^{(\ell)\top}\text{SM}\big{(}\mathbf{x}_{\theta}^{(\ell)}(\mathbf{z}_{t},t)\big{)}\cdot\mathbf{1}[\tilde{\mathbf{x}}_{0}\text{ is Invalid}]$ (13)
@@
where $\mathbf{1}[\tilde{\mathbf{x}}_{0}\text{ is Invalid}]$ denotes an indicator function that returns 1 when the sampled sequence $\tilde{\mathbf{x}}_{0}$ is invalid and 0 if valid and $k=\arg\max_{j}(\mathbf{x}^{(\ell)}_{\theta}\big{(}\mathbf{z}_{t},t)\big{)}$ is the token with the highest predicted probability at position $\ell$ of the sequence.
Differentiating the invalid loss with respect to the probability vector $\mathbf{x}_{\theta}^{(\ell)}(\mathbf{z}_{t},t)$ for position $\ell$, we derive the gradient with respect to the predicted probability of the sampled token $j=r$ and all other tokens in the vocabulary $j\neq r$ in Appendix B.1 as
@@
Minimizing this objective function updates the parameters to lower the predicted probabilities for tokens that resulted in invalid peptide SMILES and increase the probabilities of the remaining tokens while maintaining a consistent relative probability distribution.
Training. To train the MDLM to accurately approximate the true reverse transition distribution $q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{0})$ of a training sample $\mathbf{x}_{0}$ for all continuous timesteps $t=1\to 0$, we train a parameterized model $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)$ that takes the current sequence $\mathbf{z}_{t}$ and returns a $|\mathcal{V}|$-dimensional vector of predicted probabilities of transitioning to each token at time $s<t$ (Algorithm 1). For each training batch $B$, we randomly sample $|B|$ values $t\in\text{Uniform}(0,1)$ and off-set each time $t$ by $\vec{\delta}=\left[0,\frac{1}{|B|},\frac{2}{|B|},\dots,\frac{|B|-1}{|B|},1\right]$ to get $\vec{t}=(\vec{t}+\vec{\delta})\text{ mod }\mathbf{1}$ to ensure the model learns to regenerate the clean sample $\mathbf{z}_{t}$ for a wide range of time steps. After applying state-dependent masking to each training sequence and obtaining the predicted probabilities $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)$ and the discrete sequence $\tilde{\mathbf{x}}_{0}$ from greedy sampling, we minimize the total loss function $\mathcal{L}$ given by
$\mathcal{L}$ $=\mathcal{L}_{\text{NELBO}}^{\infty}+\mathcal{L}_{\text{invalid}}$ (16)
$=\frac{1}{|B|}\sum_{\mathbf{z}_{t}\in B}\bigg{(}\frac{1}{t}\sum_{\ell:\mathbf{x}^{\ell}\neq\mathbf{b}}-\log\langle\mathbf{x}_{\theta}(\mathbf{z}^{\ell}_{t}),\mathbf{x}^{\ell}\rangle+\frac{w}{t}\sum_{\ell:\mathbf{x}^{\ell}=\mathbf{b}}-\log\langle\mathbf{x}_{\theta}(\mathbf{z}^{\ell}_{t}),\mathbf{x}^{\ell}\rangle$ (17)
@@
By increasing batch size and applying dynamic batching as described in (Methods 4.1, we obtain a tighter ELBO of the true distribution $\log p(\mathbf{x}_{0})$. The model used to generate the validation results in this manuscript is trained on our in-house 8$\times$×A6000 Nvidia GPUs (50G memory) for 1600 GPU hours using the AdamW optimizer with a learning rate of 0.0003 and weight decay of 0.075.
@@
$\mathbf{z}_{s}$ $\sim\arg\max\left(\frac{\log p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})}{\mathbf{G}}\right)$ ($\mathbf{G}\in\mathbb{R}^{|\mathcal{V}|}$)
$G_{i}$ $=-\log(-\log(u_{i}+\epsilon)+\epsilon)$ ($u_{i}\sim\text{Uniform}(0,1)$)
where $G_{i}$ is the Gumbel noise applied to token $i$ and $\epsilon=1e-10$. Then, we return the newly sampled tokens only when $\mathbf{z}_{t}=\mathbf{m}$, while keeping all unmasked tokens unchanged. After $T$ timesteps, we obtain a fully unmasked sequence $\mathbf{x}$.

### v2

@@ 3 Discussion
In this work, we introduce PepTune, a generative framework that achieves multi-objective optimization directly in discrete sequence space. By leveraging MCTS for guidance, PepTune identifies Pareto-optimal peptide SMILES sequences conditioned on diverse therapeutic properties such as binding affinity, solubility, membrane permeability, and hemolysis. Unlike previous guidance methods, which struggle with gradient estimation [27] or rely on projections between continuous and discrete spaces [87], PepTune operates natively in the discrete latent space. Our approach combines exploration through batched unmasking and reward-based exploitation of classifier predictions, ensuring valid peptide structures with a state-dependent masking schedule and an invalidity loss. Most importantly, unlike recent binder design methods [12, 14, 15, 88], PepTune requires no obligate target three-dimensional structures or predicted structures (only the target sequence), enabling peptide design to conformationally diverse proteins, optimized for properties beyond local geometric interactions.
Despite its strengths, PepTune relies on synthetic peptide data (e.g., CycloPs [89]) and rare non-natural amino acids (nAAs) [40], which may increase synthesis complexity and costs. While we address this through feature-rich embeddings from a pre-trained chemical language model [45, 90], improving high-quality labeled datasets remains critical for enhancing property prediction accuracy. Furthermore, while we evaluate binding using an external, state-of-the-art docking strategy via AutoDock Vina [48], there is a lack of biophysical models for other properties optimized via PepTune, including peptide solubility, hemolysis, and membrane permeability, outside of existing predictor algorithms. As such, we are currently conducting in vitro assays to confirm these properties of our generated peptides and will update the manuscript as these results are obtained.
@@ 4 Methods
Dynamic Batching. We applied dynamic batching to handle variable-length token sequences and increase computational efficiency. Inspired by ESM-2’s dynamic batching technique [44], input SMILES are sorted by length to maximize the utility of GPU memory. The maximum tokens per GPU was set to 16,000.
SMILES Tokenization. To enable the novel generation of non-natural amino acids containing cyclizations and diverse backbone and side-chain modifications, we trained our generative diffusion model on Simplified Molecular-Input Line-Entry System (SMILES) [32] representations of peptides. We experimented with several tokenization schemes that capture common motifs in the training data to enhance the generation of valid peptide SMILES. We find that the SMILES Pair Encoding (SPE) tokenization scheme [96] with the PeptideCLM [45] vocabulary of 581 SMILES tokens and 5 special tokens with an average length of four characters per token, demonstrated superior performance, generating precise but valid peptides (Appendix D.2).
Classifier Data and Training. We trained our membrane permeability XGBoost regression model using experimentally-validated peptide SMILES consisting of 22,040 SMILES sequences obtained from the ChEMBL database [99] and 7451 sequences from the CycPeptMPDB database [95].
We collected binding affinity, solubility, hemolysis, and non-fouling data for classifier training from the PepLand and PeptideBERT datasets [39, 30]. The binding affinity data was split within groups of peptides with weak, medium, and strong binding scores with a 0.8/0.2 ratio. The classifier was trained with cross-attention between ESM-2 protein embeddings and PeptideCLM SMILES representations [47]. The hyperparameters were chosen with 50 trials of OPTUNA search [100]. For other classifiers, data were randomly shuffled and split into 0.8/0.1/0.1 ratio for train, validation, and test. XGBoost classifiers [101, 102] were applied on PeptideCLM embeddings with 50 trials of OPTUNA search for the optimal boosted tree hyperparameters.
@@ 4.2 Unconditional Masked Discrete Diffusion Model
Notation. Let $\mathbf{x}_{0}\in\{0,1\}^{|\mathcal{V}|}$ represent the one-hot vector of a token in a sequence in the training data and $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)\in\Delta^{|\mathcal{V}|}$ be the vector of predicted token probabilities across the vocabulary $\mathcal{V}$ given the current state $\mathbf{z}_{t}$ at time $t$. In most contexts, $\mathbf{x}_{0}$ will be used to denote a single token, but when discussing the full sequence, $\mathbf{x}_{0}^{(\ell)}$ is used to denote the token at position $\ell$ in the sequence. Let $T$ denote the total number of time steps in the discrete forward and reverse diffusion processes. In the context of all time steps, we expand $t$ to $t(n)\in(0,1]$ when denoting a single time step in the forward and backward diffusion process with $n\in[1\dots T-1]$. Let $s(n)=t(n)-\frac{1}{T}$ denote the previous time step in the forward process. Then, let $\mathbf{z}_{t(n)}$ and $\mathbf{z}_{s(n)}$ denote the state of a specific token at time $t(n)$ and $s(n)$ in the diffusion process, respectively. Let $\alpha_{t}(\mathbf{x}_{0}):\mathbb{R}^{|\mathcal{V}|}\to\mathbb{R}$ denote a function that takes the unmasked token $\mathbf{x}_{0}$ and outputs a value in [0,1] representing the probability of remaining unmasked at time $t$ in the forward diffusion process. Let $\mathbf{b}\in\mathbb{R}^{|\mathcal{V}|}$ denote a vector with ones at indices of peptide bond tokens and zeroes at all remaining indices and let $\mathbf{x}_{0}=\mathbf{b}$ indicate that $\mathbf{x}_{0}$ is a peptide-bond token.
State-Dependent Masking Schedule. Since all peptides follow a distinct SMILES structure consisting of un-modified or modified peptide bonds before and after each central carbon atom with an amino acid side chain, we hypothesized that applying state-dependent masking and unmasking schedules would allow the reverse diffusion process to learn to unmask the crucial structural components of a peptide SMILES that are common across all peptides before filling in the segments in-between with diverse amino acid side-chains.
Therefore, we devised a masking schedule where the probability of masking a token within a peptide bond increases at a slower rate in earlier times $t$ in the masking process compared to non-peptide bond tokens. To achieve this, we define the discrete-time log-linear masking schedule $\sigma(t)=-\log(1-t)$ for non-peptide bond tokens and the log-polynomial masking schedule $\sigma(t)=-\log(1-t^{w})$ for peptide-bond tokens. We show in Appendix B.1 that the continuous-time probability of remaining unmasked at time $t$ in the forward diffusion process is given by the function $\alpha_{t}(\mathbf{x}_{0}):\mathbb{R}^{|\mathcal{V}|}\to\mathbb{R}$ that takes the vector encoding the token $\mathbf{x}_{0}$ and returns a probability.
@@
where $\mathbf{b}$ represents the with ones at indices of peptide bond tokens and zeroes at all remaining indices, including tokens within modified peptide-bonds identified with our BondMask function (Appendix 7). Since the probability of transitioning to a [MASK] token at time $t$ is given by $1-\alpha_{t}(\mathbf{x}_{0})$, there is a lower probability $t^{w}$ for $t\in(0,1]$ of masking a peptide bond token than the probability $t$ of masking a non-peptide bond token, especially in earlier time steps for smaller $t$ (Fig. 15A). As $t\to 1$, $\alpha_{t}(\mathbf{x}_{0})\to 0$ and the probability of masking for both peptide and non-peptide bond tokens approaches 1, ensuring that the model can learn to reconstruct the full token sequence during the reverse diffusion process.
With our state-dependent masking rate $\alpha_{t}(\mathbf{x}_{0})$, we define the state-dependent forward transition matrix as
@@
To estimate the reverse posterior, we define a parameterized RoFormer model $\mathbf{x}_{\theta}(\mathbf{z}_{t},t):\mathcal{V}^{L}\times[0,1]\to\Delta^{|\mathcal{V}|}$ that takes the partially masked sequence at time $t$ and predicts a vector of token probabilities over the $|\mathcal{V}|$-dimensional simplex for each position in the sequence. By substituting $\mathbf{x}_{0}=\mathbf{x}_{\theta}(\mathbf{z}_{t},t)$ into the true reverse transition, we get the predicted reverse transition distribution.
@@
SUBS Parametrization. Following Sahoo et al. [22], we parameterize the reverse diffusion model using SUBS parametrization, which enforces zero-masking probability and carry-over unmasking. This strategy enforces the constraints applied in the forward diffusion process and has been shown to minimize perplexity [104].
@@
    Zero Masking Probability. The forward process operates under the assumption that a token can only be masked once across times $t=0\to 1$. It follows that the probability of a sequence being masked in the reverse diffusion process is zero. To enforce this property, we explicitly set the log-probability of the sequence being masked in the reverse transition distribution to zero such that if a token is unmasked at time $t$, it remains unmasked for all timesteps from $t\to 0$.
@@
    Carry-Over Unmasking. For each transition in the forward pass, all tokens either remain unchanged or are masked. Therefore, in the reverse process, when a token is unmasked at time $t$, the unmasked token is copied over all time steps from $t\to 0$. We enforce this by setting the logits vector equal to the one-hot encoding for the unmasked token.
    $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)=\mathbf{z}_{t}\;\text{ if }\;\mathbf{z}_{t}\neq\mathbf{m}$ (10)
@@ 4.3 Loss Functions
Training on samples masked for continuous values of $t\sim\text{Uniform}(0,1)$ yields a tighter lower bound compared to discrete values of $t$ [105]. When the predicted probability distribution $\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)$ is exactly the one-hot encoding vector $\mathbf{x}_{0}^{(\ell)}$ for each token at position $\ell$ in the true sequence, the loss reduces to 0, which supports our objective.
@@
$\mathcal{L}^{\infty}_{\text{NELBO}}=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\mathbb{E}_{q(\mathbf{z}_{t}|\mathbf{x}_{0})}\bigg{[}-\sum_{\ell:\mathbf{x}^{(\ell)}_{0}=\mathbf{b}}\frac{w}{t}\log\langle\mathbf{x}^{(\ell)}_{0},\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\rangle-\sum_{\ell:\mathbf{x}^{(\ell)}_{0}\neq\mathbf{b}}\frac{1}{t}\log\langle\mathbf{x}^{(\ell)}_{0},\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\rangle\bigg{]}$ (12)
@@
Given that peptide bonds form the fundamental backbone structure of a peptide, our state-dependent masking strategy for peptide bonds acts as a peptide bond loss that introduces a higher penalty when the token predictions at positions of peptide bonds are inconsistent from the ground truth tokens during training, forcing the model to learn the specific structure of peptide SMILES strings in a vast space of SMILES strings that are not valid peptides.
Invalid Peptide Loss. To further discourage the generation from predicting token logits that produce invalid peptide SMILES, we incorporate a loss to penalize sampling of invalid peptide SMILES during training by taking the argmax of the predicted logits and assigning a penalty based on our peptide validity filter (Appendix 8). Given the peptide sequence $\tilde{\mathbf{x}}^{(\ell)}\in\{0,1\}^{K}$ generated from the argmax tokens with the highest probability from the predicted logits $\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)$, we minimize a penalty determined by our validity filter $\mathbf{1}[\tilde{\mathbf{x}}_{0}\text{ is Invalid}]$ which returns 0 when the sequence is a valid peptide SMILES and 1 when the sequence either is not a valid SMILES or does not correspond to a peptide sequence. Since the argmax function is not differentiable, we use the softmax probability of the sampled tokens to scale the penalty score for each token which acts as a scalar multiplier in the loss function.
$\mathcal{L}_{\text{invalid}}$ $=\sum_{\ell=1}^{L}\tilde{\mathbf{x}}_{0}^{(\ell)\top}\text{SM}\big{(}\mathbf{x}_{\theta}^{(\ell)}(\mathbf{z}_{t},t)\big{)}\cdot\mathbf{1}[\tilde{\mathbf{x}}_{0}\text{ is Invalid}]$
@@
where $k=\arg\max_{j}(\mathbf{x}^{(\ell)}_{\theta}\big{(}\mathbf{z}_{t},t)\big{)}$ is the token with the highest predicted probability at position $\ell$ of the sequence.
Differentiating the invalid loss with respect to the probability vector $\mathbf{x}_{\theta}^{(\ell)}(\mathbf{z}_{t},t)$ for position $\ell$, we derive the gradient with respect to the predicted probability of the sampled token $j=k$ and all other tokens in the vocabulary $j\neq k$ in Appendix B.1 as
@@
Minimizing this objective function updates the parameters to lower the predicted probabilities for tokens that result in invalid peptide SMILES and increase the probabilities of the remaining tokens proportional to their original distribution such that the relative probability distribution of all other tokens $j\neq k$ is maintained.
Training. To train the MDLM to accurately approximate the true reverse transition distribution $q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{0})$ of a training sample $\mathbf{x}_{0}$ for all continuous timesteps $t=1\to 0$, we train a parameterized model $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)$ that takes the current sequence $\mathbf{z}_{t}$ and returns a $|\mathcal{V}|$-dimensional vector of predicted probabilities of transitioning to each token at time $s<t$ (Algorithm 1). For each dynamic training batch $B$, we randomly sample $|B|$ values $t\in\text{Uniform}(0,1)$ and off-set each time $t$ by $\vec{\delta}=\left[0,\frac{1}{|B|},\frac{2}{|B|},\dots,\frac{|B|-1}{|B|},1\right]$ to get a vector $\vec{t}=(\vec{t}+\vec{\delta})\text{ mod }\mathbf{1}$ of evenly distributed time steps to ensure the model learns to regenerate the clean sample $\mathbf{z}_{t}$ for a continuous range of time steps. After applying state-dependent masking to each training sequence and obtaining the predicted probabilities $\mathbf{x}_{\theta}(\mathbf{z}_{t},t)$ and the discrete sequence $\tilde{\mathbf{x}}_{0}$ from greedy argmax sampling, we minimize the total loss function $\mathcal{L}$ given by
$\mathcal{L}$ $=\mathcal{L}_{\text{NELBO}}^{\infty}+\mathcal{L}_{\text{invalid}}$
$=\frac{1}{|B|}\sum_{\mathbf{x}_{0}\in B}\bigg{(}-\sum_{\ell:\mathbf{x}^{(\ell)}_{0}=\mathbf{b}}\frac{w}{t}\log\langle\mathbf{x}^{(\ell)}_{0},\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\rangle-\sum_{\ell:\mathbf{x}^{(\ell)}_{0}\neq\mathbf{b}}\frac{1}{t}\log\langle\mathbf{x}^{(\ell)}_{0},\mathbf{x}^{(\ell)}_{\theta}(\mathbf{z}_{t},t)\rangle$ (15)
@@
By increasing batch size and applying dynamic batching (Methods 4.1), we obtain a tighter ELBO of the true distribution $\log p(\mathbf{x}_{0})$. The model used to generate the validation results in this manuscript is trained on our in-house 8$\times$×A6000 Nvidia GPUs (50G memory) for 1600 GPU hours using the AdamW optimizer with a learning rate of 0.0003 and weight decay of 0.075.
@@
$\mathbf{z}^{(\ell)}_{s}$ $\sim\arg\max\left(\log p_{\theta}(\mathbf{z}^{(\ell)}_{s}|\mathbf{z}^{(\ell)}_{t})+\mathbf{G^{(\ell)}}\right)$ ($\mathbf{G}^{(\ell)}\in\mathbb{R}^{|\mathcal{V}|}$)
$G^{(\ell)}_{i}$ $=-\log(-\log(u^{(\ell)}_{i}+\epsilon)+\epsilon)$ ($u^{(\ell)}_{i}\sim\text{Uniform}(0,1)$)
where $G_{i}$ is the Gumbel noise applied to the token probability at index $i$ for position $\ell$ and $\epsilon=1e-10$. Then, we return the newly sampled tokens only when $\mathbf{z}^{(\ell)}_{t}=\mathbf{m}$, while keeping all unmasked tokens unchanged. After $T$ timesteps, we obtain a fully unmasked sequence $\mathbf{x}$.